OpenStudy (anonymous):

if a cylindrical barrel measures 22 inches in diameter, how many inches will it roll in 8 revolutions along a smooth surface?

OpenStudy (anonymous):

Well, how many inches do you think it will roll in 1 revolution?

OpenStudy (anonymous):

4 inches?

OpenStudy (anonymous):

Draw a picture, draw more than one. This about what is rolling. Think about what is making contact with the ground. A small hint, the circumference is somehow involved. Bedtime now. good luck

OpenStudy (anonymous):

In one revolution the barrel will roll a distance of Pi*22 inches the length of a circle's circumference. In eight revolutions 8*Pi*22 = 552.93 feet.

OpenStudy (anonymous):

The circumference of the cylindric barrel will tell us how many inches it rolls in 1 revolution. The circumference of a circle is Πd. Since the diameter is 22: \[Π d * 22\] This equals.. 22/7 x 22 = 69.14 inches per revolution. For 8 revolutions, multiply by 8. \[69.14 \times 8 = 553.12\] The cylindrical barrel travels 553.12 inches in 8 revolutions.